8                     Given that pH = - log [H+]


(a)              Find the pH of a solution which has a [H+] concentration of 


                                               i.     0.01 mol dm-3


                                             ii.     0.004 mol dm-3


            Determine the [H+] concentration of a solution with a pH of 3.8


1
Expert's answer
2021-03-09T06:09:18-0500

(a)

i.

"pH=-\log(0.01)=2"

ii.


"pH=-\log(0.004)=2.398"

(b)


"pH=-\log[H^+]=>[H^+]=10^{-pH}"

"[H^+]=10^{-3.8}=0.0001585\ mol\cdot dm^{-3}"